OpenStudy (anonymous):

2. Solve the system by substitution. 2x - y + z = -4 z= 5 -2x +3y - z = -10 (1 point) (–8, 7, 5) (–8, –7, 5) (8, –7, 5) (–8, –7, –5)

OpenStudy (anonymous):

To begin, try substituting z=5 into the first equation and solve for y.

OpenStudy (anonymous):

Substituting z=5 into the first equation, you get: 2x-y+z=-4 2x-y+(5)=-4 2x-y=-9 2x+9=y So, y=2x+9. Now, substitute y=2x+9 into the third equation.

OpenStudy (anonymous):

Substituting z+5 and y=2x+9 into the third equation, you get: -2x+3y-z=-10 -2x+3(2x+9)-(5)=-10 -2x+6x+27-5=-10 4x+22=-10 4x=-32 x=-8 So, you now know that x=-8 and z=5. Substitute x=-8 into y=2x+9 to find y.

OpenStudy (anonymous):

okay. Thank you so much. I was having trouble understanding this problem but now I get it.

OpenStudy (anonymous):

You're very welcome.

OpenStudy (anonymous):

does Y= 25 or did i solve it wrong?

OpenStudy (anonymous):

Remember that x=-8, so when you substitute it into y=2x+9 you get: y=2x+9 y=2(-8)+9 y=-16+9 y=-7

OpenStudy (anonymous):

oh okay. thanks again.

OpenStudy (anonymous):

Yeah. No problem.